OpenStudy (anonymous):

Write an equation for the table. x -1 0 1 2 y 5 3 1 -1 Answer y = -2x + 3 y = -x + 3 y = 2x + 5 y = x + 5

OpenStudy (kc_kennylau):

constant is 3 (refer to x=0)

OpenStudy (kc_kennylau):

y=mx+3 substitute (1,1): 1=m+3 m=-2 \(\therefore y=-2x+3\rightarrow A\)